# Rambo: RL-augmented Model-based Whole-body Control for Loco-manipulation

### v2

@@ IV-B Generating Feedforward Torque via WBC
In addition to the desired acceleration ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{\bm{a}}$, Rambo also accept user’s specification of desired EE forces ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{\bm{u}}_{i}$ for the manipulation end-effectors $i\in\mathcal{M}$, as shown in Figure 4. Using ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\hat{\bm{a}}$ and ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\hat{\bm{u}}_{i}$ in the base frame ${}\{\mathcal{B}\}$, we formulate the following QP to optimize the reaction force
$\min_{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\bm{u}_{i},i\in\mathcal{N}}\left\lVert\Delta{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}{\bm{a}}\right\rVert_{\bm{U}}^{2}+\sum_{i\in\mathcal{M}}\left\lVert\Delta{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\bm{u}_{i}\right\rVert_{\bm{V}}^{2}+\sum_{i\in\mathcal{L}}\left\lVert{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\bm{u}_{i}\right\rVert_{\bm{W}}^{2}$ (10a)
$\text{subject}\ \text{to}:{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}{\bm{a}}$ $=\sum_{i=1}^{N}\bm{A}_{i}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\bm{u}_{i}+\hat{\bm{g}}$ (10b)
$({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\bm{u}_{i})_{z}$ $=0,$ $i\in\mathcal{L}_{\text{swing}}$ (10c)
$\left\lVert({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\bm{u}_{i})_{x}\right\rVert$ $\leq\mu({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\bm{u}_{i})_{z},$ $i\in\mathcal{L}_{\text{stance}}$ (10d)
$\left\lVert({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\bm{u}_{i})_{y}\right\rVert$ $\leq\mu({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\bm{u}_{i})_{z},$ $i\in\mathcal{L}_{\text{stance}}$ (10e)
$u_{z,\min}$ $\leq({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\bm{u}_{i})_{z}\leq u_{z,\max},\quad$ $i\in\mathcal{L}_{\text{stance}}$ (10f)
@@
where $\Delta{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}{\bm{a}}={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\hat{\bm{a}}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}{\bm{a}}$, $\Delta{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\bm{u}_{i}={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\hat{\bm{u}}_{i}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\bm{u}_{i}$; ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}{\bm{a}}=[{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}{\dot{\bm{v}}},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}{\dot{\bm{\omega}}}]\in\mathbb{R}^{6}$ is the acceleration of the single rigid body; $\bm{A}_{i}$ is the generalized inverse inertia matrix defined in Equation 4; $\mu$ is the friction coefficient; $\mathcal{L}_{\text{swing}},\mathcal{L}_{\text{stance}}$ are the set of end-effectors for locomotion in swing and stance respectively; $\bm{U},\bm{V},\bm{W}\succ 0$ are positive definite weight matrices. We note that friction checking is only performed on locomotion end-effectors due to the uncertainty of contact surface for manipulation. Leveraging the SRB model, Rambo efficiently accounts for the dynamic effects from contacts.
@@
$\bm{\tau}^{\text{FF}}_{j}=\sum_{i=1}^{N}\bm{J}_{i}^{\top}\cdot{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\bm{u}_{i},$ (11)
@@ IV-C Learned Policy
We design the observation space $\mathcal{O}$ to include the proprioceptive information, gait information, kinematic joint position target, and user commands. They are chosen to ensure reward function can be successfully induced from only the observation. We stack 6-step history of observations as input to the policy [44]. In contrast to previous works [42, 43], the action $a_{t}\in\mathcal{A}$ is designed to have two separate heads: base acceleration correction $\Delta{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\hat{\bm{a}}}$ and joint position correction $\Delta\hat{\bm{q}}_{j}$, providing feedback to both feedforward torque calculation and joint positions. The surrogate targets are
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\tilde{\bm{a}}}$ $={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\hat{\bm{a}}}+\Delta{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\hat{\bm{a}}}$ (14)
@@
where ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\tilde{\bm{a}}}$ is taken as the target for the base acceleration for the reaction force optimization module. The desired joint position $\tilde{\bm{q}}_{j}$ is used to calculate the final joint torque command sent to the robot, formulated as
@@
Projected gravity ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\bm{g}$ 3
Base linear velocity ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\bm{v}$ 3
Base angular velocity ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\bm{\omega}$ 3
@@
Velocity command $({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{v}_{x},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{v}_{y},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{\omega}_{z})$ 3
EE position command ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\hat{\bm{r}}^{m}_{i}}$ $3\cdot|\mathcal{M}|$
EE force command ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{\bm{u}}_{i}$ $3\cdot|\mathcal{M}|$
@@
TABLE I: Detailed observation space $\mathcal{O}$. $\bm{q}_{j0}$ is the default joint position; $\bm{\phi}\in[-1,1]$ is the phase for each periodic limb motion; $\bm{\psi}\in\{-1,0,1\}$ is the mode for each limb to distinguish the periodic patterns (swing, gait, stance); ||$\cdot$|| ⋅ | is the cardinality of the set.
@@
Base orientation ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\bm{g}_{z}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\hat{\bm{g}}_{z}$ 0.3 / 0.6
Linear velocity ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\bm{v}}_{x,y}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{\bm{v}}_{x,y}$ 0.2 / 0.3
Angular velocity ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\bm{\omega}}_{z}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{\bm{\omega}}_{z}$ 0.3 / 0.4
EE position ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{{\bm{r}}^{m}_{i}}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\hat{\bm{r}}^{m}_{i}}$ 0.1
$r_{\text{reg}}$ Contact mismatch $\hat{c}_{i}\neq c_{i}$
@@
TABLE II: Detailed reward functions. To ensure all reward functions produce values with the range of $[0,1]$, we map the error term for each reward using $r=\exp(\left\lVert\bm{\epsilon}\right\rVert^{2}/\sigma^{2})$. The contact mismatch reward is mapped using $r=0.5^{|\hat{c}_{i}\neq c_{i}|}$. ()/()($\cdot$)/($\cdot$)( ⋅ ) / ( ⋅ ) indicates different values for quadruped or biped tasks. We set desired base height $\hat{\bm{p}}_{z}$ to 0.3 and 0.45, desired gravity vector ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\hat{\bm{g}}_{z}$ to $[0,0,-1]$ and $[-1,0,0]$ for quadruped and biped tasks.

### v4

@@ IV-B Generating Feedforward Torque via WBC
In addition to the desired acceleration ${{}_{\mathcal{P}}}\hat{\bm{a}}$, Rambo also accept user’s specification of desired EE forces ${{}_{\mathcal{P}}}\hat{\bm{u}}_{i}$ for the manipulation end-effectors $i\in\mathcal{M}$, as shown in Figure 4. Using $^{{}_{\mathcal{B}}}\hat{\bm{a}}$ and ${{}_{\mathcal{B}}}\hat{\bm{u}}_{i}$ in the base frame $\{\mathcal{B}\}$, we formulate the following QP to optimize the reaction force
$\min_{{{}_{\mathcal{B}}}\bm{u}_{i},i\in\mathcal{N}}\left\lVert\Delta{{}_{\mathcal{B}}}{\bm{a}}\right\rVert_{\bm{U}}^{2}+\sum_{i\in\mathcal{M}}\left\lVert\Delta{{}_{\mathcal{B}}}\bm{u}_{i}\right\rVert_{\bm{V}}^{2}+\sum_{i\in\mathcal{L}}\left\lVert{{}_{\mathcal{B}}}\bm{u}_{i}\right\rVert_{\bm{W}}^{2}$ (10a)
$\text{subject}\ \text{to}:{{}_{\mathcal{B}}}{\bm{a}}$ $=\sum_{i=1}^{N}\bm{A}_{i}{{}_{\mathcal{B}}}\bm{u}_{i}+\hat{\bm{g}}$ (10b)
$({{}_{\mathcal{B}}}\bm{u}_{i})_{z}$ $=0,$ $i\in\mathcal{L}_{\text{swing}}$ (10c)
$\left\lVert({{}_{\mathcal{B}}}\bm{u}_{i})_{x}\right\rVert$ $\leq\mu({{}_{\mathcal{B}}}\bm{u}_{i})_{z},$ $i\in\mathcal{L}_{\text{stance}}$ (10d)
$\left\lVert({{}_{\mathcal{B}}}\bm{u}_{i})_{y}\right\rVert$ $\leq\mu({{}_{\mathcal{B}}}\bm{u}_{i})_{z},$ $i\in\mathcal{L}_{\text{stance}}$ (10e)
$u_{z,\min}$ $\leq({{}_{\mathcal{B}}}\bm{u}_{i})_{z}\leq u_{z,\max},\quad$ $i\in\mathcal{L}_{\text{stance}}$ (10f)
@@
where $\Delta{{}_{\mathcal{B}}}{\bm{a}}={{}_{\mathcal{B}}}\hat{\bm{a}}-{{}_{\mathcal{B}}}{\bm{a}}$, $\Delta{{}_{\mathcal{B}}}\bm{u}_{i}={{}_{\mathcal{B}}}\hat{\bm{u}}_{i}-{{}_{\mathcal{B}}}\bm{u}_{i}$; ${{}_{\mathcal{B}}}{\bm{a}}=[{{}_{\mathcal{B}}}{\dot{\bm{v}}},{{}_{\mathcal{B}}}{\dot{\bm{\omega}}}]\in\mathbb{R}^{6}$ is the acceleration of the single rigid body; $\bm{A}_{i}$ is the generalized inverse inertia matrix defined in Equation 4; $\mu$ is the friction coefficient; $\mathcal{L}_{\text{swing}},\mathcal{L}_{\text{stance}}$ are the set of end-effectors for locomotion in swing and stance respectively; $\bm{U},\bm{V},\bm{W}\succ 0$ are positive definite weight matrices. We note that friction checking is only performed on locomotion end-effectors due to the uncertainty of contact surface for manipulation. Leveraging the SRB model, Rambo efficiently accounts for the dynamic effects from contacts.
@@
$\bm{\tau}^{\text{FF}}_{j}=\sum_{i=1}^{N}\bm{J}_{i}^{\top}\cdot{{}_{\mathcal{B}}}\bm{u}_{i},$ (11)
@@ IV-C Learned Policy
We design the observation space $\mathcal{O}$ to include the proprioceptive information, gait information, kinematic joint position target, and user commands. They are chosen to ensure reward function can be successfully induced from only the observation. We stack 6-step history of observations as input to the policy [44]. In contrast to previous works [42, 43], the action $a_{t}\in\mathcal{A}$ is designed to have two separate heads: base acceleration correction $\Delta{{}_{\mathcal{P}}}{\hat{\bm{a}}}$ and joint position correction $\Delta\hat{\bm{q}}_{j}$, providing feedback to both feedforward torque calculation and joint positions. The surrogate targets are
${{}_{\mathcal{P}}}{\tilde{\bm{a}}}$ $={{}_{\mathcal{P}}}{\hat{\bm{a}}}+\Delta{{}_{\mathcal{P}}}{\hat{\bm{a}}}$ (14)
@@
where ${{}_{\mathcal{P}}}{\tilde{\bm{a}}}$ is taken as the target for the base acceleration for the reaction force optimization module. The desired joint position $\tilde{\bm{q}}_{j}$ is used to calculate the final joint torque command sent to the robot, formulated as
@@
Projected gravity ${{}_{\mathcal{B}}}\bm{g}$ 3
Base linear velocity ${{}_{\mathcal{B}}}\bm{v}$ 3
Base angular velocity ${{}_{\mathcal{B}}}\bm{\omega}$ 3
@@
Velocity command $({{}_{\mathcal{P}}}\hat{v}_{x},{{}_{\mathcal{P}}}\hat{v}_{y},{{}_{\mathcal{P}}}\hat{\omega}_{z})$ 3
EE position command ${{}_{\mathcal{P}}}{\hat{\bm{r}}^{m}_{i}}$ $3\cdot|\mathcal{M}|$
EE force command ${{}_{\mathcal{P}}}\hat{\bm{u}}_{i}$ $3\cdot|\mathcal{M}|$
@@
TABLE I: Detailed observation space $\mathcal{O}$. $\bm{q}_{j0}$ is the default joint position; $\bm{\phi}\in[-1,1]$ is the phase for each periodic limb motion; $\bm{\psi}\in\{-1,0,1\}$ is the mode for each limb to distinguish the periodic patterns (swing, gait, stance); $|\cdot|$ is the cardinality of the set.
@@
Base orientation ${{}_{\mathcal{B}}}\bm{g}_{z}-{{}_{\mathcal{B}}}\hat{\bm{g}}_{z}$ 0.3 / 0.6
Linear velocity ${{}_{\mathcal{P}}}{\bm{v}}_{x,y}-{{}_{\mathcal{P}}}\hat{\bm{v}}_{x,y}$ 0.2 / 0.3
Angular velocity ${{}_{\mathcal{P}}}{\bm{\omega}}_{z}-{{}_{\mathcal{P}}}\hat{\bm{\omega}}_{z}$ 0.3 / 0.4
EE position ${{}_{\mathcal{P}}}{{\bm{r}}^{m}_{i}}-{{}_{\mathcal{P}}}{\hat{\bm{r}}^{m}_{i}}$ 0.1
$r_{\text{reg}}$ Contact mismatch $\hat{c}_{i}\neq c_{i}$ -
@@
TABLE II: Detailed reward functions. To ensure all reward functions produce values with the range of $[0,1]$, we map the error term for each reward using $r=\exp(\left\lVert\bm{\epsilon}\right\rVert^{2}/\sigma^{2})$. The contact mismatch reward is mapped using $r=0.5^{|\hat{c}_{i}\neq c_{i}|}$. $(\cdot)/(\cdot)$ indicates different values for quadruped or biped tasks. We set desired base height $\hat{\bm{p}}_{z}$ to 0.3 and 0.45, desired gravity vector ${{}_{\mathcal{B}}}\hat{\bm{g}}_{z}$ to [0,0,1][0,0,-1] and [1,0,0][-1,0,0] for quadruped and biped tasks.